ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 62 K 99, 62 K 05 Naumov, Anatoly, Novosibirsk State Technical University, Russia Design of Experiments and its regularization The investigations, carried by us, had testified that system active identification problem in general and experimenting strategy synthesis problem in particular is incorrectly formulated ones, and the using of Kiefer-Wolfowitz(J. Kiefer,1959,1974) statements in practice leads to falling down of efficiency of experimental plans. It may be said, that the active identification problem in a classic form is a twice incorrectly raised one. Regrettably, the BoxDraper’s ( J. E. P. Box, N. R. Draper,1965) approach doesn’t solve all the matters of regularization of incorrectly - raised active identification problems, and in some cases does reinforce this incorrectness.